# Designing Parallel Algorithms for Community Detection using Arachne

### v1

@@ III-B2 Aggregation Phase
In the implementation of the aggregation phase of the Louvain algorithm, we adopt a streamlined but robust strategy to remap and construct higher-level graphs iteratively. After determining the updated community assignments for each vertex at the start of the aggregation phase, the community IDs are first remapped to a contiguous range starting from 0 to the total number of unique communities. This ensures efficient indexing and facilitates following graph processing.
Subsequently, the edge list is traversed to update the endpoints of each edge according to the new community IDs. During this step, edge weights are recalculated based on the interactions between the communities represented by the endpoints. This updated edge list, now representing the coarsened graph, is transmitted back to the Python front end.
The Arachne framework leverages Arkouda’s GroupBy and Broadcast functionality during the graph reconstruction process. This critical step automatically merges edges with identical endpoints, aggregating their weights to produce the final edge weights for the coarsened graph. This integration of GroupBy and Broadcast not only simplifies the graph reconstruction but also ensures computational efficiency by offloading the aggregation process to Arkouda’s optimized backend.
Following this, the algorithm alternates between the local-moving and aggregation phases in an iterative manner. During each iteration, vertices are reassigned to communities in the local-moving phase based on modularity optimization, and the edge list is updated in the aggregation phase to reflect the new community assignments and then reconstruct a new graph. These steps are repeated until no further changes in vertex community assignments occur, representing convergence and the termination of the algorithm. The overall process is outlined in the following pseudo-code.
@@
6:           $done\leftarrow|\{c\mid c\in C\}|==|V(G)|$triangleright Convergence
7:           if not $done$ then
@@
14:     $V\leftarrow C$triangleright Remapping
@@ IV-B Experimental configuration
The experiments were conducted on an AMD EPYC 7763 server with dual sockets and 512GB of memory. Each socket contains 64 cores, providing a total of 128 cores. The system is equipped with 64KB of L1 caches, 512KB L2 caches, and 32MB L3 caches. Arachne is set up to work in the client-server model, where the client, typically a Python script or Jupyter notebook, interacts with an Arkouda server running on a high-performance computing (HPC) system. The relevant code was compiled using GCC version 13.2.0.
@@ V-A Comparing performance
We evaluate the performance of our parallel implementations of two community detection algorithms: Label Propagation and Louvain using Arachne. These implementations are compared to established Python libraries, including igraph, NetworkX and NetworKit. The runtime for each algorithm is measured five times, and the average is reported to ensure accuracy and consistency.
@@ V-A1 Label Propagation Algorithm
As shown in Fig 1, our experimental results demonstrate that Arachne-LPA exhibits superior performance across all test graphs. For smaller graphs such as com-dblp and com-amazon, Arachne-LPA achieves execution times of 0.5s and 0.4s respectively, demonstrating speedups of 96.9x and 42.7x compared to igraph implementations. This performance advantage becomes more pronounced with larger graphs: on com-livejournal, Arachne-LPA completes in 16.3s compared to igraph’s 425.6s (26.2x speedup), while on com-orkut, it achieves a 15.4x speedup over igraph (31.9s vs 492.5s). Notably, NetworkX-LPA shows significant performance degradation on larger graphs, requiring 3686.4s for com-orkut.
Figure 1: Runtime in seconds with igraph-LPA, NetworkX-LPA, and Arachne-LPA
@@ V-A2 Louvain Algorithm
The runtime comparison of different Louvain implementations is presented in Fig. 2. The Arachne-Louvain implementation consistently outperforms other methods across almost all tested datasets. For smaller graphs like com-dblp and com-amazon, Arachne completes in 1.54s and 1.57s, respectively, compared to 5.50s and 4.00s for igraph, 41.75s and 30.33s for NetworkX, and 0.44s and 0.62s for NetworKit. While NetworKit achieves lower runtimes on some small graphs, Arachne demonstrates strong scalability and becomes increasingly advantageous on larger datasets. For instance,on com-livejournal Arachne completes in 5.59s, significantly faster than igraph (417.34s), NetworkX (2214.39s), and NetworKit (69.31s). The largest performance gain appears on com-orkut, where Arachne completes in 15.60s, compared to 1176.58s for igraph, 11072.52s for NetworkX, and 109.14s for NetworKit.
Figure 2: Runtime in seconds with igraph-Louvain, NetworkX-Louvain, NetworKit-Louvain, and Arachne-Louvain
As shown in Fig. 3, Arachne-Louvain achieves modularity that are highly comparable to those produced by sequential implementations. For example, on com-amazon and as-skitter, Arachne reaches modularity values of 0.9248 and 0.8289, closely matching the results from igraph (0.9262 and 0.8404) and NetworkX (0.9260 and 0.8462). In contrast, NetworKit tends to produce lower modularity values on several datasets, such as 0.6918 on com-amazon and 0.7503 on as-skitter. These results suggest that Arachne not only provides significantly faster execution than sequential implementations but also delivers higher quality results than existing parallel baselines like NetworKit.
@@ V-A3 Analysis across Graph Sizes
The experimental results reveal consistent performance characteristics across the two algorithms. Our experiments demonstrate that the relative performance advantage of Arachne implementations exhibits a strong correlation with graph size. For smaller graphs such com-dblp and com-amazon, we observe moderate but consistent performance improvements. However, the efficiency becomes substantially more significant for large-scale graphs like com-livejournal and com-orkut, where performance improvements frequently span multiple orders of magnitude. This scaling behavior suggests that Arachne’s parallel processing approach becomes particularly effective as graph size increases, making it especially suitable for large-scale graph analysis tasks.
@@ V-B Strong Scaling
The strong scaling experiments were conducted on the com-livejournal dataset to evaluate the parallel efficiency of our Louvain implementation, as shown in Fig.4. We varied the number of threads from 2 to 128 by configuring the CHPL_RT_NUM_THREADS_PER_LOCALE environment variable in Chapel. To provide deeper insights into the scaling behavior, we also break down the runtime of different phases to better understand the individual performance.
Figure 4: Scalability of execution times as the number of threads is increased for the Louvain implementation.
@@
When it comes to phase-level analysis, the local-moving phase, which dominates the sequential execution time in Louvain algorithm, shows the most significant improvement with parallelization. Specifically, its execution time improves by 7x when scaling from 2 to 128 threads, reducing from 16.8s to 2.4s. This indicates that the label-update operations in this phase are highly parallelizable. In contrast, the aggregation phase exhibits limited scalability due to its global communication requirements.
@@ VI Related Work
Numerous community detection algorithms have been developed, tackling the problem from various perspectives.
Label propagation is a widely used local heuristic for community detection. Introduced by Raghavan et al.[5], it gained popularity due to its simplicity, efficiency, and ease of implementation. Various improvements have since been proposed to address its limitations. For example, Zhang[23] introduced edge clustering coefficients into the update rule to enhance stability and accuracy, while Wang [24] proposed NI-LPA, which incorporates a vertex importance index—computed via k-core decomposition, neighborhood signal propagation, and Jaccard similarity—into the label selection process. These enhancements aim to improve the robustness and quality of label propagation methods.
The Louvain algorithm is a greedy modularity optimization method for detecting communities with high modularity. To scale it to large graphs, several parallel Louvain methods (PLMs) have been proposed. A key challenge in PLMs is the synchronization of vertex updates. Some methods use non-real-time strategies where vertex states are updated only after each full iteration. For instance, Lu [9] implemented a non-real-time Louvain variant using OpenMP. In contrast, real-time strategies update vertex information immediately during local moves. Naim [25] adopts this approach using two hash tables to manage vertex and community data, achieving up to 270× speedup on medium-scale graphs. However, such methods can incur high storage and computation costs.
Other work focuses on leveraging hardware acceleration. Forster [26] employs CUDA to implement Louvain on GPUs, achieving up to 12× speedup compared to CPU-based versions. Bhowmick et al.[27] proposed HyDetect, a hybrid CPU-GPU Louvain algorithm, and later extended it to support multi-node, multi-GPU systems[28]. Chou’s Nido system applies a batched clustering strategy optimized for GPUs, enabling processing of graphs exceeding aggregate GPU memory.
Several graph analytics libraries provide community detection capabilities. NetworKit [29] is a hybrid C++/Python framework offering parallel implementations of Louvain and Leiden [30]. cuGraph, part of NVIDIA’s RAPIDS suite, delivers GPU-accelerated graph analytics through a Python interface, with core routines written in CUDA/C++.
@@ VII Conclusion and Future Work
Large graph analysis presents significant challenges for data scientists, particularly in community detection, which is a common task. In this paper, we have developed and implemented several parallel and optimized algorithms for community detection with the open-source framework, Arachne. Arachne facilitates large-scale graph analytics by leveraging supercomputers and cloud resources while providing a user-friendly Python interface. Our experimental results show that our implementation within the Arachne framework achieves substantial performance improvements, with absolute speedups compared to existing Python toolkits like NetworkX, igraph, and NetworKit, demonstrating great scalability as well.
For future work, we plan to expand the capabilities of Arachne by implementing additional community detection algorithms. In the meantime, we aim to conduct comparative experiments against other state-of-the-art implementations to further validate the performance advantages of Arachne. By continuing to enhance and optimize our software, we try to empower data scientists to analyze and derive insights from large-scale graph data.
@@ VIII Acknowledgement
We thank the Chapel and Arkouda communities for their support and guidance. This research is supported in part by NSF grants CCF-2109988, OAC-2402560. and CCF-2453324.

### v2

@@ III-B2 Aggregation Phase
In the implementation of the aggregation phase of the Louvain algorithm, we adopt a streamlined but robust strategy to iteratively remap and construct higher-level graphs. After determining the updated community assignments for each vertex at the start of the aggregation phase, the community IDs are first remapped to a contiguous range from 0 to the total number of unique communities. This ensures efficient indexing and facilitates the subsequent graph processing.
Subsequently, the edge list is traversed to update the endpoints of each edge according to the new community IDs. During this step, edge weights are recalculated based on the interactions between the communities represented by the endpoints. This updated edge list, now representing the coarsened graph, is transmitted back to the Python front-end.
The Arachne framework leverages Arkouda’s GroupBy and Broadcast functionality during the graph reconstruction process. This critical step automatically merges edges with identical endpoints, aggregating their weights to produce the final edge weights for the coarsened graph. This integration of GroupBy and Broadcast not only simplifies the graph reconstruction, but also ensures computational efficiency by offloading the aggregation process to Arkouda’s optimized back-end.
Following this, the algorithm alternates between the local-moving and aggregation phases in an iterative manner. During each iteration, the vertices are reassigned to communities in the local-moving phase based on modularity optimization, and the edge list is updated in the aggregation phase to reflect the new community assignments and then reconstruct a new graph. These steps are repeated until no further changes in vertex community assignments occur, representing convergence and the termination of the algorithm. The overall process is outlined in the following pseudo-code.
@@
6:      $done\leftarrow|\{c\mid c\in C\}|==|V(G)|$ $\triangleright$ Convergence
7:      if not donedone then
@@
14:  $V\leftarrow C$ $\triangleright$ Remapping
@@ IV-B Experimental configuration
The experiments were conducted on an AMD EPYC 7763 server with dual sockets and 512GB of memory. Each socket contains 64 cores, providing a total of 128 cores. The system is equipped with 64KB L1 caches, 512KB L2 caches, and 32MB L3 caches. Arachne is set up to work in the client-server model, where the client, typically a Python script or Jupyter notebook, interacts with an Arkouda server running on a high-performance computing (HPC) system. The relevant code was compiled using GCC version 13.2.0. We used NetworKit v11.0, NetworkX v3.1, and igraph v0.11.8 as baselines. Unless otherwise stated, all running times exclude I/O and graph loading time and only include the algorithmic execution.
@@ V-A Comparing performance
We evaluate the performance of our parallel implementations of two community detection algorithms: Label propagation and Louvain, using Arachne. These implementations are compared with established Python libraries, including igraph, NetworkX, and NetworKit. The running time and modularity for each algorithm are measured five times, and the average is reported to ensure accuracy and consistency.
@@ V-A1 Label Propagation Algorithm
As shown in Fig. 1, our experimental results demonstrate that Arachne-LPA exhibits superior performance on all test graphs. For smaller graphs such as com-dblp and com-amazon, Arachne-LPA achieves execution times of 0.5s and 0.4s, respectively, demonstrating speedups of 96.9x and 42.7x compared to igraph implementations. This performance advantage becomes more pronounced with larger graphs: on com-livejournal, Arachne-LPA completes in 16.3s compared to igraph’s 425.6s (26.2x speedup), while on com-orkut, it achieves a 15.4x speedup over igraph (31.9s vs 492.5s). In particular, NetworkX-LPA shows significant performance degradation on larger graphs, requiring 3686.4s for com-orkut.
Figure 1: Running time in seconds with igraph-LPA, NetworkX-LPA, and Arachne-LPA
@@ V-A2 Louvain Algorithm
The comparison of the running time of different Louvain implementations is presented in Fig. 2. The Arachne-Louvain implementation consistently outperforms other methods in almost all the datasets tested. For smaller graphs like com-dblp and com-amazon, Arachne completes in 1.54s and 1.57s, respectively, compared to 5.50s and 4.00s for igraph, 41.75s and 30.33s for NetworkX, and 0.44s and 0.62s for NetworKit. Although NetworKit achieves lower running times on some small graphs, Arachne demonstrates strong scalability and becomes increasingly advantageous on larger datasets. For example, on com-livejournal, Arachne completes in 5.59s, significantly faster than igraph (417.34s), NetworkX (2214.39s), and NetworKit (69.31s). The greatest performance gain appears on com-orkut, where Arachne completes in 15.60s, compared to 1176.58s for igraph, 11072.52s for NetworkX, and 109.14s for NetworKit.
Figure 2: Running time in seconds with igraph-Louvain, NetworkX-Louvain, NetworKit-Louvain, and Arachne-Louvain
As shown in Fig. 3, Arachne-Louvain achieves modularity that is highly comparable to those produced by sequential implementations. For example, on com-amazon and as-skitter, Arachne reaches modularity values of 0.9248 and 0.8289, closely matching the results from igraph (0.9262 and 0.8404) and NetworkX (0.9260 and 0.8462). In contrast, NetworKit tends to produce lower modularity values on several datasets, such as 0.6918 on com-amazon and 0.7503 on as-skitter. These results suggest that Arachne not only provides significantly faster execution than sequential implementations but also delivers higher-quality results than existing parallel baselines like NetworKit.
@@ V-A3 Analysis across Graph Sizes
The experimental results reveal consistent performance characteristics between the two algorithms. Our experiments demonstrate that the relative performance advantage of Arachne implementations exhibits a strong correlation with the size of the graph. For smaller graphs such as com-dblp and com-amazon, we observe moderate but consistent performance improvements. However, efficiency becomes substantially more significant for large-scale graphs like com-livejournal and com-orkut, where performance improvements often span multiple orders of magnitude. This scaling behavior suggests that Arachne’s parallel processing approach becomes particularly effective as graph size increases, making it especially suitable for large-scale graph analysis tasks.
@@ V-B Strong Scaling
Strong scaling experiments were conducted on the com-livejournal dataset to evaluate the parallel efficiency of our Louvain implementation, as shown in Fig.4. We varied the number of threads from 2 to 128 by configuring the CHPL_RT_NUM_THREADS_PER_LOCALE environment variable in Chapel. To provide deeper insights into the scaling behavior, we also break down the running time of different phases to better understand the individual performance.
Figure 4: Scalability of execution times as the number of threads is increased for the Louvain implementation
@@
When it comes to phase-level analysis, the local-moving phase, which dominates the sequential execution time in the Louvain algorithm, shows the most significant improvement with parallelization. Specifically, its execution time improves by 7x when scaling from 2 to 128 threads, reducing from 16.8s to 2.4s. This indicates that the label-update operations in this phase are highly parallelizable. In contrast, the aggregation phase exhibits limited scalability due to its global communication requirements.
@@ VI Related Work
Community detection has been extensively studied to address the problem from various perspectives, with label propagation [5] and Louvain algorithm [6] being two of the most widely used heuristics.
Label propagation has gained popularity due to its simplicity, efficiency, and ease of implementation. Various improvements have since been proposed to address its limitations [23, 24]. The Louvain algorithm is a greedy modularity optimization method for detecting communities with high modularity. To scale it to large graphs, several parallel Louvain methods (PLMs) have been proposed [9, 25]. They demonstrated that parallel heuristics can achieve modularity comparable to sequential baselines while scaling to much larger graphs.
Several graph analysis frameworks expose these algorithms through Python interfaces. NetworkX [26] is widely adopted in the Python ecosystem, but its pure Python implementation lacks parallelization. igraph [27] provides a C back-end with Python bindings, offering improved performance but only limited multithreading. NetworKit [28] is a high-performance C++ framework with Python bindings that includes parallel Louvain (PLM) and label propagation (PLP) implementations. Our comparisons focus on these frameworks because they represent the most widely used Python-facing tools, and thus provide the most relevant baselines to evaluate Arachne’s parallel performance.
Other state-of-the-art implementations, such as GPU-based algorithms [29, 30] or distributed-memory algorithms [8], target different execution environments. Although these achieve excellent performance on their respective platforms, they are beyond the scope of this work.
@@ VII Conclusion and Future Work
Large graph analysis presents significant challenges for data scientists, particularly in community detection, which is a common task. In this paper, we have developed and implemented several parallel and optimized algorithms for community detection with the open-source framework Arachne. Arachne facilitates large-scale graph analytics by leveraging supercomputers and cloud resources while providing a user-friendly Python interface. Our experimental results show that our implementation within the Arachne framework achieves substantial performance improvements, with absolute speedups compared to existing Python toolkits such as NetworkX, igraph, and NetworKit, demonstrating great scalability as well.
For future work, we plan to expand the capabilities of Arachne by implementing additional community detection algorithms. In the meantime, our aim is to conduct comparative experiments against other state-of-the-art implementations to further validate the performance advantages of Arachne. By continuing to improve and optimize our software, we try to empower data scientists to analyze and derive insights from large-scale graph data.
@@ VIII Acknowledgement
We thank the Chapel and Arkouda communities for their support and guidance. This research is supported in part by NSF grants CCF-2109988, OAC-2402560, and CCF-2453324.